OpenStudy (anonymous):

True or False. The equation -2m + 5 = 2m + 5 has a solution set : ∅

OpenStudy (sweetburger):

How many values of m satisfy the equation?

OpenStudy (mathmale):

Why not actually try solving the equation first?

OpenStudy (ethane):

i want to say t

OpenStudy (ethane):

r

OpenStudy (ethane):

u

OpenStudy (ethane):

e

OpenStudy (mathmale):

I'd rather that you do some work before saying "true" or "false." Given -2m + 5 = 2m + 5, please ADD 2m to both sides of the equation.

OpenStudy (anonymous):

I just wanna know what that symbol means, I know how to solve or simplify the given equation.

OpenStudy (ethane):

what symbol?

OpenStudy (anonymous):

the O with a line.

OpenStudy (anonymous):

How do you know if that is what a 'solution set' equals?

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

-2m + 5 = 2m + 5 Adding 2m to both sides results in 5=4m+5. True or false?

OpenStudy (anonymous):

True, I want to know what the O with a line symbol means.

OpenStudy (mathmale):

"empty set"

OpenStudy (anonymous):

So what would that look like?

OpenStudy (mathmale):

This would mean "NO solution." However, I ask you to finish solving the problem. Pls go back to my last equation. Can you solve for m?

OpenStudy (mathmale):

This is an empty set: { }. It has squiggly brackets, denoting "set," but there's nothing inside the brackets!

OpenStudy (anonymous):

Why, yes I can. You then subtract 5 from both sides, resulting in 4m.

OpenStudy (mathmale):

Yes, You'd have 4m on the right side. But what would you have on the left side?

OpenStudy (anonymous):

0

OpenStudy (mathmale):

Yes, and therefore 4m=0. Solve this for m.

OpenStudy (anonymous):

You divide 4 from both sides so is it an empty set?

OpenStudy (anonymous):

Or multiply by the recipricol lol

OpenStudy (anonymous):

Don't know why I said lol but yea

OpenStudy (mathmale):

4m=0 dividing this by 4 results in \[\frac{ 4m }{ 4 }=\frac{ 0 }{ 4 }, ~or~m=?\]

OpenStudy (anonymous):

m=0 so....

OpenStudy (mathmale):

Yes. There is a solution. Your solution set is {0}. This is NOT an empty set. An empty set would look like { }.

OpenStudy (anonymous):

What is an example of an empty set?

OpenStudy (mathmale):

{ } is an empty set: it has nothing in it. An empty jar would be regarded as an empty set. Nothing there!

OpenStudy (anonymous):

I meant an equation. Nevermind then, I guess. Thank you.

OpenStudy (mathmale):

Nice working with you. Take care, Sheyenne.

OpenStudy (anonymous):

I agree, you too.